OpenStudy (anonymous):

Compare and contrast, in complete sentences, the difference between the products of y2y5 and (y2)5. Provide the expanded form and product of each.

OpenStudy (anonymous):

@celecity

OpenStudy (anonymous):

Ah, okay, so, what do you know about the rules of exponents?

OpenStudy (anonymous):

I think its like this y^2y^5 = y^7 Because you add the exponents

OpenStudy (anonymous):

Yep, right! Have any ideas for the second one?

OpenStudy (anonymous):

but with the parenthese you'd multiply

OpenStudy (anonymous):

Precisely.

OpenStudy (anonymous):

(y^2)^5 = y^10 Because you multiply the exponents.

OpenStudy (anonymous):

Yep. So as far as the difference in product go, they're going to be different because adding and multiplying are going to give you a different number. Let's say y equals 2. (2)^2*(2)^5 2^7 128 (2^2)^5 2^10 1024 The difference is in the ^7 vs. the ^10. Does that makes sense?

OpenStudy (anonymous):

I see!

OpenStudy (anonymous):

The expanded form is going to look like this: (y^2)(y^5) yyyyyyy (y^2)^5 (y^2)(y^2)(y^2)(y^2)(y^2)

OpenStudy (anonymous):

So comparing the two, they both have exponents of 2 and 5. The difference is, the first one you add, and the second on you multiply

OpenStudy (anonymous):

The product of the first one would be y^7 and the product of the second one would be y^10

OpenStudy (anonymous):

thanks so much celecity

OpenStudy (anonymous):

You're very welcome :)

OpenStudy (anonymous):

:)